Rosa has designed a pulley system for a school science project. She has determined that as the coefficient of friction increases, the efficiency of the pulley decreases. With a coefficient of friction of 0.2, the efficiency of the pulley is 75%. What would you expect the efficiency to be if the coefficient of friction increased to 0.3?

Answer by stanbon(75887)

Rosa has designed a pulley system for a school science project. She has determined that as the coefficient of friction increases, the efficiency of the pulley decreases.
e = k/f
-----------------
Solve for "k" using "With a coefficient of friction of 0.2, the efficiency of the pulley is 75%".
0.75 = k/0.2
k = 0.75*0.2 = 0.15
--------------------
Equation:
e = 0.15/f
----
What would you expect the efficiency to be if the coefficient of friction increased to 0.3?
e = 0.15/0.30
==============
Cheers,
Stan H.
--------------
coefficient of friction = 1/2